Mal87:

If f(x)=x^2 and g(x)=x+3, which of the following statements is true? A.(f*g)(0)=3 B.(g*f)(1)=16 C.(f*g)(3)=36 D. (f*g)(1)=2

Komaeda:

hm. I think it's supposed to be f composed with g not f times g that being said, we can start by finding f composed with g by taking the function g(x) = x + 3 and substituting that into the x value of f(x) so f ยบ g(x) = g(x)^2 = (x+3)^2 so start plugging in the x-values for choices A, C, and D to see which one produces the given result